La Silla Paranal Observatory News

The page contains important news and events of the La Silla Paranal Observatory. It also covers some important telescope and instruments news from the La Silla, Paranal and APEX sites.

General News

Paranal News

Telescope News


  • Nov 20, 2013: recoating of M1, M3 of UT2 completed
  • Sept 11, 2013: recoating of M1, M3 of UT4 completed
  • April 1, 2013: The VST is currently out of service between March 29, 2013, and April 2, 2013 due to a multiple damage of cables and cooling lines after an incident caused by the failure of a limit switch. All damaged parts were replaced or repaired.
  • February 26, 2013: The Laser Guide Star system at UT4 has recently been successfully upgraded with a fibre Raman laser (AOF prototype) delivering improved on-sky performances as well as enhanced operational stability and maintenance. Full characterization of the upgraded system will soon be available. This intervention will result in a dramatic increased in the availability of the LGS system for science operations.


Instrument News


  • December 21: VISIR: sensitivities in imaging and low resolution spectroscopy have been updated for phase 2 of P97. Also an M-band filter has been partially commissioned. See the VISIR news page.
  • December 21: Start of a SINFONI upgrade: it includes replacement of the J, H, K and H+K filters, replceament of the pre-optics with better J-band transmission, replacement of mirrors with better surface quality ones to improve the line spread function in particular in the J-band. Re-commissioning is expected to take place on January 29 and 30.
  • December 15: Publications of early science results of the XXL survey, which aimed at providing a well-defined sample of about 500 galaxy clusters out to a distance when the Universe was half of its current age: five new superclusters of galaxies, fewer distance clusters than expected, etc... The survey combines the XMM-Newton survey and with data obtained from ground based observatories, including from EFOSC2, FORS and MUSE.
  • December 14: NACO is back in operation. All SM modes are operational, but some VM modes need some additional work. The camera wheel has recovered its functionality. Quadrant 4 has also been recovered; however, quadrant 2 (lower left) has not been recovered: 1 in 8 columns has no signal. See the NACO news.
  • December 9: Observations with FORS and MUSE reveals the aftermath of a 360 million year collision between galaxies, including a peculiar young dwarf galaxy.
  • December 2-17: A problem with an electronic board of SAXO impedes observations with SPHERE. See the SPHERE news.
  • November 28: A 6.2 earthquake hit Paranal on November 27 at 18:00 local time: its epicenter was very close (within 30 km) to the observatory. After the necessary checks, most instruments were shown to work nominally afterwards. However, the UVES blue arm camera shows coma which reduces its spectral resolution.
  • November 25: SPHERE obtained the most detailed image of the hypergiant star VY Canis Majoris.
  • November 25: MUSE observations lead to several predictions regarding the appearance of the Refsdal supernova lensed by a galaxy cluster.
  • November 23: First PIONIER service mode observations with the upgraded VLTI.
  • November 20 - December 4: Maintenance of the VIRCAM cooling system  and software upgrade.
  • November 18: The UltraVISTA survey uncovered 574 new massive galaxies but none earlier than about obe billion years after the Big Bang.
  • November 16: The Call for proposals for the VISIR Upgrade Science Verification has been issued. See the VISIR news page.
  • November 11: Doppler tomography using several instruments, including UVES and X-shooter, over a twelve-years period details the shattered remains of an asteroid around a white dwarf.
  • November 4-17: first, successful, commissioning run of GRAVITY.
  • October 28: The VISTA Variables  in the Via Lactea survey discovers a disc of young stars buried behind thick coulds in the Mily Way central Bulge.
  • October 21: FLAMES finds hottest and most massive touching double star.
  • October 7: Comparison between images obtained with SPHERE and previous NASA/ESA Hubble images revealed fast-moving wave-like features in the disc of the star AU Mic.
  • October 1: Start of a major intervention on NACO to recover the camera wheel and improve the detector characteristics.
  • September 4: The AG slide of X-shooter will have to be repaired in the coming months. See the X-shooter news.
  • September 1: The first commissioning of GRAAL, part of the Adaptive Optics Facility, is completed. HAWK-I is back in operation.
  • August 25: First fringes with the VLTI following the  upgrade: MIDI as been removed from the laboratory, PIONIER now sits above FINITRO, the laboratory can now accomodate GRAVITY and MATISSE, new relay optic systems have been installed on the ATs, a new AT maintenance station has been commissioned, new hardware has been installed for GRAVITY, new delay line relay mirroes have been commissioned for new operation stations, the UT Coudé rooms have been prepared for the GRAVITY wavefron sensors (CIAO)
  • August 14-26: Unstability of the temperature of both detectors lead to suspend observations with VISIR. See VISIR news page.
  • August 12: Twilight flats obtained after the FORS2 LADC prism exchange have now been analysed, confirming that the structure clearly visible in the old data is gone. See the FORS2 News page.
  • August 10: The Galaxy and Mass Assembly (GAMA) survey releases first data, obtained with VIRCAM and OmegaCAM, as well as NASA GALEX and WISE and ESA Herschel. It meausures the energy output of over 200,000 galaxies across as broad a wavelength range as possible.
  • August 1: A Call for Letters of Intent for new public surveys with VISTA was released. See the VIRCAM News page.
  • July 28: FLAMES: Spectral resolutions have improved following the update in the focussing procedure: see the FLAMES News page.
  • July 19: MUSE: detector voltage configuration files for Channels 1 and 9 were changed. See the MUSE news page.
  • July 9: First release of KiDS, the dark matter survey using the OmegaCAM.
  • July 8: X-shooter spectra show that some long-duration gamma-ray bursts are driven by magnetars.
  • July 7: VISIR is back in operation, as the problem with its astatic support has been fixed.
  • June 29: NACO is back in operation, with the camera wheel fixed on S13. However, despite 8 attempts, it was not possible to recover the detector fully: NACO will thus be operated with 3 quadrants only.
  • June 25: FLAMES data provides evidence that M87 accreted a ~ 6 x 109 Msun galaxy in the last giga-year. 
  • June 23: VIRCAM is back in operation. A short power cut today affecting all instruments caused further problems but were quickly solved this time.
  • June 19: KMOS is back in operation.
  • June 18: UVES is back in operation following the power cut. Problems still affect KMOS. If the detector is fine, NACO could be back in operation on June 19th. VIRCAM should be operational at the earliest on June 21.
  • June 17: X-shooter and SINFONI helped to provide the best observational evidence of first generation stars in the Universe.
  • June 17: HAWK-I is recovered.
  • June 16: FORS2 is back in operation.
  • June 15: FLAMES-GIRAFFE, SPHERE, VIMOS, MUSE, and OmegaCAM are back in operation following the power cut.
  • June 14: Passively cooled instruments could be back in operation as fast as the rest of the infrastructure allowed, such as X-shooter and SINFONI. However problems still affect UT1, UT3, VST and many systems.
  • June 12: A power cut affected the whole observatory: caused by the failure of a back-up generator while the main (gas) turbine was undergoing maintenance. No operation took place on the nights of June 12th and 13th.
  • June 10: SPHERE reveals earliest stage of planetary nebula formation.
  • May 21: VISIR is warming up in advance of an intervention aimed at fixing a problem with the astatic support found as the cause of the observed drift. It is expected that the instrument would be back in operation on July 7th.
  • May 13: FLAMES discovers a new kind of globular star cluster
  • May 12: The cause of the drift seen with VISIR has been identified and requires a new intervention to fix the problem. 
  • April 30: MUSE reveals The Pillars of Creation in 3D
  • April 16: Together with Hubble, SINFONI observations show that star formation shuts down in the centres of elliptical galaxies first
  • April 15: Has MUSE found the First Signs of Self-interacting Dark Matter?
  • April 2: SPHERE enters regular operation. Note however that most DIT/NDITs need to be updated. See the SPHERE News page for details.
  • April 2:  A drift problem affects observations for most targets with VISIR, limiting objects to bright ones for imaging and excluding spectroscopic observations.
  • March 26: SINFONI has provided the Best View Yet of Dusty Cloud Passing Galactic Centre Black Hole
  • March 20: NACO: The mechanism that allows to switch objectives suffered a severe failure and CONICA had to be warmed up to be serviced. See more details in the NACO News.
  • March 11: VST- OmegaCAM finds A Grand Extravaganza of New Stars
  • March 5: Start of the work to refurbush the VLTI Laboratory. MIDI has been decommissioned. Note that the instrument web page has been moved! The VLTI will be offered again in Period 96.
  • March 5: CRIRES has mapped the distribution of H2O and HDO on Mars: The Planet that Lost an Ocean’s Worth of Water
  • March 04: KMOS: After an intervention to upgrade all arms, the instrument is handed back to Science Operation for recommissioning and testing.
  • February 26: MUSE goes beyond Hubble: Looking Deeply into the Universe in 3D
  • February 18: First SPHERE Science results: The Strange Case of the Missing Dwarf
  • Feburary 9: FLAMES:
    • The cleaning of the Giraffe HR grating was again successfully performed at the end of January 2015, resulting in an increase in efficiency estimated to be in the range 10-20% for bluer settings.
    • On the afternoon of February 2, 2015, an update of the focussing procedure was implemented in FLAMES/Giraffe. This results in significant increase in both spectral resolution and sensitivity across all settings, although these enhencements vary from setting to setting. See the FLAMES News page.
  • February 9: Stellar Partnership Doomed to End in Catastrophe
  • February 4: VISTA Stares Right Through the Milky Way
  • January 22: KMOS Arm 2 was put in simulation.
  • January 17: NACO is resuming operation. The new Aladdin2 (former ISAAC LW) detector installed in CONICA features slightly different noise properties than the Aladdin3. In addition, there is a significant additional thermal background that will slightly reduce the dynamic range and detection limits when long exposure times are used, associated with short wavelength broadband filters. See more details in the NACO News.


  • December 4: SPHERE Science Verification has started.
  • December 2: KMOS  Arm 19 was put into simulation
  • November 21: KMOS  Arm 8 was put into simulation.
  • November 18: KMOS  Arm 4 was put into simulation.
  • November 15: The FORS2 LADC, which was suffering from a poor coating, was replaced by the one of FORS1 whose coating had been removed. This led to a gain in throughput of 0.1 ( in the blue), 0.07 (in V), 0.05 (in I), and 0.04 (in I) magnitudes, due to a reduction of the scattering. See the FORS2 news page for updates.
  • October 6: KMOS Arms 5,6,7,and 16 have been replaced, Electronics of Arm4 were changed. All arms except 11 are back in operation.
  • NACO: updates following the move of NACO to UT1 can be seen on the NACO news page.
  • August 24: KMOS arm 7 is parked, locked and in simulation
  • August 6, 7: KMOS arm 5 is parked, locked and in simulation; arm 15 is recovered and back in operation
  • July 17: last night of operation for CRIRES before the start of the upgrade
  • July 7: on July 4, an anomalous noise was heard coming from inside the OMEGACAM's cryostat. By precaution, the instrument was out of operation on the nights of July 4 and 5th.
  • June 20 - 29: MUSE Science Verification
  • June 4: press release describing the first light of SPHERE.
  • June 2: speeds of the grating wheel of SINFONI had to be decreased
  • May 31: KMOS arm 15 is parked, locked and in simulation
  • May 20: KMOS arms 4 and 11 are temporarily out of operations.
  • May 20: FLAMES ETC with improved efficiency measures at all settings.
  • May 10: start of the SPHERE commissioning.
  • April 1: KMOS Arm 11 is back in operation.
  • March 5: press release describing the first light of MUSE.
  • March 1: Limiting magnitudes for AMBER in the K-band has increased, owing to polarization control.
  • February 21: KMOS was successfully recommissioned after an intervention to lower the vibration and to repair the arms that were locked during the last months. Arm 11 is still not available, all other arms work well within the specifications.
  • February 7: start of the MUSE commissioning.
  • January 7: KMOS arm 11 is parked, locked and in simulation.
  • January 7: UVES red arm CCD shutter failure. See the UVES News page.


La Silla News

The section provides the latest telescope and instruments news from the La Silla Site.

More news can be found in the news section of theLa Silla Science Operations page.

  • March 14, 2010: Updated service observing information Service Observing
  • November 20, 2009: A new technique using near-infrared images, obtained with s 3.58-metre New Technology Telescope (NTT)ESO Watching a Cannibal Galaxy Dine
  • April 6, 2009: La Silla Astroclimatology information (weather, seeing, extinction) was moved to a new location and can now be found on the mean navigation menu of the La Silla Facilities pages. Additional La Silla astroclimatology information will be added here.
  • February 11, 2009: From the second week of July, 2009, observations will be carried out from a new control room located in theadministration building.
  • February 11, 2009: A new page describing the way that operations will run after P83 (ie from October 2009) is available here.
  • April 22, 2008: A document that explains how to interface visitor instruments with focal stations of La Silla telescopes is available (2.8 MB PDF format).
  • March 27, 2007: The dome of the 3.6m has been fully restored, and moves with nominal speed.
  • October 05, 2006: The dome of the 3.6m enters an extended period of corrective maintenance. Science operations at the 3.6m is temporarily suspended until further notice.
  • May 17, 2006: A Trio of Neptunes  discovered with HARPS, see ESO PR 18/06
  • January 26, 2006: A 5.5M_earth planet discovered with the 1.54m Danish telescope.